# On solutions of singular Sylvester equations in quaternions

### 24 Mar 2026 v3

@@ 3. Quaternion square roots
For an arbitrary nonzero quaternion a{a} (3.1) has the solutions
@@
The square roots of an arbitrary nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
@@
The square roots of an arbitrary nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
@@
For an arbitrary nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
@@ 4. Solutions to the homogeneous singular Sylvester equation
is referred to as a homogeneous singular Sylvester equation. In this section, we restate the existance and the general solution to (4.1) previously presented in [6]. Moving beyond existing theory, we provide a novel characterization of the nontrivial solutions to (4.1) by applying the quaternion square root, presented in Theorem 4.3.
@@
${x}={a}{q}+{q}{a}=2{a}\cdot(\mathrm{Re}\,{q}{}-{\mathrm{Im}\,{q}}).$
Let $\lambda,\mu$ be arbitrary real numbers. Then, with ${a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b})$ and ${a}\cdot{\mathrm{Im}\,{q}}=\lambda\sqrt{{a}{b}^{*}}$, according to Lemma 3.2, we obtain (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
@@
Hence, x{x} is a pure quaternion that satisfies ${a}\cdot{x}=2{a}\cdot({a}\times{\mathrm{Im}\,{q}})=0$. Moreover, by Lemma 3.2, and with (a+b)=0({a}+{b})=0, we obtain (4.4). Across all cases, the solution in (4.4) is nonzero if and only if $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu({a}+{b})\mathclose{\rvert}\neq 0$, which completes the proof. ∎
@@ 5. Solutions to the inhomogeneous singular Sylvester equation
If a,b{a},{b} are real numbers, (5.1) has no solution, since (ab)x=0({a}-{b}){x}=0 for all quaternions x{x}. Therefore, in the following theorems, we assume that a,b{a},{b} are nonreal quaternions. Analogous to Lemma 2.7, we investigate the solutions to (5.1) for the following casses. Case 1: ${a}{b}\neq{b}{a}$, Case 2: ab=ba{a}{b}={b}{a}. Then, by Lemma 2.1 it holds either Case 2a: a=b{a}={b} or Case 2b: ${a}={b}^{*}$. The solvability condition influences the structure of the solution.
@@
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$. Then, (5.1) has a solution if and only if ${a}{c}={c}{b}^{*}$.
@@
Corollary 5.1.1.
Case 2a: Let a=b{a}={b}. Then, (5.1) has a solution if and only if $({\mathrm{Im}\,{a}}){c}=-{c}({\mathrm{Im}\,{a}})$. Case 2b: Let ${a}={b}^{*}$. Then, (5.1) has a solution if and only if ac=ca{a}{c}={c}{a}.
@@
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for an arbitrary quaternion q{q}, the general solution to (5.1) is given as
@@
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for arbitrary real numbers $\lambda,\mu$ the general solution to (5.1) is given as
@@
From Lemma 4.2 and Theorem 4.3 for an arbitrary quaternion q{q} and arbitrary real numbers $\lambda,\mu$ it holds

### 27 Mar 2026 v4

@@ 3. Quaternion square roots
For a nonzero quaternion a{a} (3.1) has the solutions
@@
The square roots of a nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
@@
The square roots of a nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
@@
For a nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
@@ 4. Solutions to the homogeneous singular Sylvester equation
is referred to as a homogeneous singular Sylvester equation. In this section, we restate the existance and the general solution to (4.1) previously presented in [6] and provide constructive proofs. Moving beyond existing theory, we also provide a novel characterization of the nonzero solutions to (4.1) by applying the quaternion square root, presented in Theorem 4.3.
@@
${x}={a}{q}+{q}{a}=2{a}(\mathrm{Re}\,{q}{})-2{a}\cdot{\mathrm{Im}\,{q}}.$
Then, with real numbers $\mu=\mathrm{Re}\,{q}{}$ and $\lambda=-2({a}\cdot{\mathrm{Im}\,{q}})/\mathopen{\lvert}{a}\mathclose{\rvert}$ it follows
$2{a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b}),\qquad-2{a}\cdot{\mathrm{Im}\,{q}}=\lambda\mathopen{\lvert}{a}\mathclose{\rvert}=\lambda\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}=\lambda\sqrt{{a}{b}^{*}},$
which yields (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
@@
Since x{x} is a pure quaternion that satisfies ${a}\cdot{x}=2{a}\cdot({a}\times{\mathrm{Im}\,{q}})=0$, with $\lambda=\mathopen{\lvert}{x}\mathclose{\rvert}/\mathopen{\lvert}{a}\mathclose{\rvert}$ and by Lemma 3.2 we have
${x}=\lambda\mathopen{\lvert}{a}\mathclose{\rvert}\frac{{x}}{\mathopen{\lvert}{x}\mathclose{\rvert}}=\lambda\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}\frac{{x}}{\mathopen{\lvert}{x}\mathclose{\rvert}}=\lambda\sqrt{-\mathopen{\lvert}{a}\mathclose{\rvert}^{2}}=\lambda\sqrt{{a}{b}^{*}},$
where the expression on the right hand side must satisfy ${a}\cdot\sqrt{{a}{b}^{*}}=0$. Then, with (a+b)=0({a}+{b})=0 we obtain (4.4). Across all cases, the solution in (4.4) is nonzero if and only if $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu({a}+{b})\mathclose{\rvert}\neq 0$, which completes the proof. ∎
@@ 5. Solutions to the inhomogeneous singular Sylvester equation
If a,b{a},{b} are real numbers, (5.1) has no solution, since (ab)x=0({a}-{b}){x}=0 for all quaternions x{x}. Therefore, in the following theorems, we assume that a,b{a},{b} are nonreal quaternions. The solvability condition influences the structure of the solution.
@@
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion. Then, (5.1) has a solution if and only if ${a}{c}={c}{b}^{*}$.
@@
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion with ${a}{c}={c}{b}^{*}$. Then, for an arbitrary quaternion q{q}, the general solution to (5.1) is given as
@@
Let a,b{a},{b} be similar quaternions and c{c} a nonzero quaternion with ${a}{c}={c}{b}^{*}$. Then, for arbitrary real numbers $\lambda,\mu$ the general solution to (5.1) is given as
@@
From Lemma 4.2 and Theorem 4.3, for an arbitrary quaternion q{q} there exist real numbers $\lambda,\mu$ such that